# One Map to Find Them All: Real-time Open-Vocabulary Mapping for Zero-shot Multi-Object Navigation

### v1

@@ II Related Work
Efforts in object goal navigation [4] aim to equip mobile systems with the ability to use geometric and semantic properties of a given environment to efficiently guide an agent towards a target object. A large body of work has studied semantic goal navigation with a predefined set of object classes (c.f. [5][6][7]). Advances in large pre-trained vision-language models have further allowed the object navigation problem to be solved for objects and environments without prior training. In several recent works, such VLM-based approaches have been used for zero-shot object navigation.
@@ III-A Problem Formulation
From a stream of RGB and depth images with associated camera poses we construct a two-dimensional, open-vocabulary belief map $\mathcal{M}=(F,\sigma^{2})$ over CLIP-aligned features $F(x,y)\in\mathbb{R}^{n_{x}\times n_{y}\times f}$ and a variance per map location $\sigma^{2}(x,y)\in\mathbb{R}^{n_{x}\times n_{y}}$, where $f$ is the CLIP feature dimension. We index the map $\mathcal{M}$ with $x\in[0,n_{x}],y\in[0,n_{y}]$ and data points in image-space with $i\in[0,H],j\in[0,W]$, where $n_{x}$, $n_{y}$ are the map’s dimensions in grid cells and $H$, $W$ are the image height and width respectively in pixels.
@@ V-A Single-Object Navigation
Our superior performance in SPL indicates that our proposed semantic feature map provides more informative guidance, resulting in shorter paths to the target object. However, it is interesting that we also outperform VLFM on success rate alone, as this does not factor in path length and we use the same object detector. Moreover, VLFM uses a PointNav locomotion policy trained specifically to better traverse the HM3D domain while we simply compute the shortest path to the frontier on a grid. We suspect that our higher SR compared to VLFM on the single-target task can be attributed to our ability to do consensus-filtering with our high-confidence semantic beliefs as outlined in Section III-D, whereas VLFM only uses heuristics to mitigate false detections. 3D reconstruction datasets such as HM3D additionally suffer from artifacts [3] which can exacerbate misdetections. To investigate our surprisingly good results on single-object SR further, we repeat the evaluation without any consensus-filtering or heuristics to mitigate false detections at all. Instead, we just accept all detections from the detector. The results of this ablation shown in Table II indicate that false positives are a considerable source of error on the ObjectNav task and our filtering can effectively reduce the false positive rate (FP) of the object detector by $7.1\%$. We note that we still achieve higher SPL than VLFM (with its misdetection heuristics and domain-specific locomotion policy). We attribute this to our semantic feature map simply being more informative. Note that our consensus-filtering requires spatially accurate features, which our map provides, as opposed to VLFM, as shown in Fig. 4, and hence would not yield the same benefit if applied to less accurate maps.
@@
Overall, we conclude that our semantic belief map can more effectively guide an agent for single-object navigation tasks, and moreover it also reduces false positives from the object detection stage.
Target Object
Agent
@@ V-B Multi-Object Navigation
We evaluate the performance of our method OneMap against VLFM on the multi-object navigation task in HM3D. Table III highlights our method’s superiority over the baseline VLFM across all metrics. We achieve an average progress (PR) of $65.54\%$, meaning that we successfully find 2 objects per episode on average. Moreover, as expected the performance gap to VLFM grows larger for multi-object tasks than for single-object tasks, indicating that our method can effectively reuse the map for multiple objects.
@@
We evaluate this further in Fig. 5, where we compute the average SPL for the first, second, and third object goals of each episode, given that the previous object was successfully found. As evident in the figure, the efficiency (SPL) for our method monotonically increases, whereas it remains relatively constant for VLFM. This can be attributed to our uncertainty-aware exploration resulting in high-quality open-vocabulary navigation maps that can effectively reuse the data from previous object goals, unlike VLFM, which builds a query-specific scoring map that cannot leverage prior searches. As a result, our method sees a relative increase in efficiency (SPL) by a factor of 1.5x compared to VLFM for the third object, benefiting from the high-quality open-vocabulary semantic map constructed during the search for the first and second objects. We note that $100\%$ SPL here represents perfect (oracle) knowledge of where every object is in advance and is therefore an unattainable upper bound.
@@ V-C Real World Experiments
We deploy our method on a Boston Dynamics Spot quadruped using only one front-facing Realsense D455 camera, and a Livox Mid 360 lidar with Fast-LIO2 [44] for odometry. We use Yolo-World [40] instead of YOLOv7 to allow searching arbitrary objects, instead of MS-COCO [39] classes only. We run the complete system on-board a Jetson Orin AGX, and achieve an overall update frequency of $2\,\mathrm{Hz}$, which we found to be sufficient for an indoor, legged robot.
@@ VI Conclusion
Although our methods works well with projection 2D, future work could include examining suitable 3D map representations, though this will introduce additional complexity to the mapping and planning parts of our method.

### v2

@@ II Related Work
Efforts in object goal navigation [4] aim to equip mobile systems with the ability to use geometric and semantic properties of a given environment to efficiently guide an agent towards a target object. A large body of work has studied semantic goal navigation with a predefined set of object classes (c.f. [5][6]). Advances in large pre-trained vision-language models have further allowed the object navigation problem to be solved for objects and environments without prior training. In several recent works, such VLM-based approaches have been used for zero-shot object navigation.
@@ III-A Problem Formulation
From a stream of RGB and depth images with associated camera poses we construct a two-dimensional, open-vocabulary belief map $\mathcal{M}=(F,\sigma^{2})$ over CLIP-aligned features $F(x,y)\in\mathbb{R}^{n_{x}\times n_{y}\times f}$ and a variance per map location $\sigma^{2}(x,y)\in\mathbb{R}^{n_{x}\times n_{y}}$, where $f$ is the CLIP feature dimension. We index the map $\mathcal{M}$ with $x\in[0,n_{x}],y\in[0,n_{y}]$ and data points in image-space with $i\in[0,H],j\in[0,W]$, where $n_{x}$, $n_{y}$ are the map’s dimensions in grid cells and $H$, $W$ are the image height and width respectively in pixels. The memory requirements grow linearly with the number of observed cells. For instance, a map of $50\,\mathrm{m}\times 50\,\mathrm{m}$ with 10 cells per meter requires roughly $700\,\mathrm{Mb}$ of memory.
@@ V-A Single-Object Navigation
Our superior performance in SPL indicates that our proposed semantic feature map provides more informative guidance, resulting in shorter paths to the target object. However, it is interesting that we also outperform VLFM on success rate alone, as this does not factor in path length and we use the same object detector. Moreover, VLFM uses a PointNav locomotion policy trained specifically to better traverse the HM3D domain while we simply compute the shortest path to the frontier on a grid. We suspect that our higher SR compared to VLFM on the single-target task can be attributed to our ability to do consensus-filtering with our high-confidence semantic beliefs as outlined in Section III-D, whereas VLFM only uses heuristics to mitigate false detections. 3D reconstruction datasets such as HM3D additionally suffer from artifacts [3] which can exacerbate misdetections. To investigate our surprisingly good results on single-object SR further, we repeat the evaluation without any consensus-filtering or heuristics to mitigate false detections at all. Instead, we just accept all detections from the detector. The results of this ablation shown in Table II indicate that false positives are a considerable source of error on the ObjectNav task and our filtering can effectively reduce the false positive rate (FP) of the object detector by $7.1\%$. We note that we still achieve higher SPL than VLFM (with its misdetection heuristics and domain-specific locomotion policy). We attribute this to our semantic feature map simply being more informative. Note that our consensus-filtering requires spatially accurate features, which our map provides, as opposed to VLFM, as shown in Fig. 4, and hence would not yield the same benefit if applied to less accurate maps. Overall, we conclude that our semantic belief map can more effectively guide an agent for single-object navigation tasks, and moreover it also reduces false positives from the object detection stage.
Target Object
Agent
@@ V-B Multi-Object Navigation
We evaluate the performance of our method OneMap against VLFM on the multi-object navigation task in HM3D. Table III highlights our method’s superiority over the baseline VLFM across all metrics. We achieve an average progress (PR) of $65.54\%$, meaning that we successfully find 2 objects per episode on average. Moreover, as expected the performance gap to VLFM grows larger for multi-object tasks than for single-object tasks, indicating that our method can effectively reuse the map for multiple objects. We evaluate this further in Fig. 5, where we compute the average SPL for the first, second, and third object goals of each episode, given that the previous object was successfully found. As evident in the figure, the efficiency (SPL) for our method monotonically increases, whereas it remains relatively constant for VLFM. This can be attributed to our uncertainty-aware exploration resulting in high-quality open-vocabulary navigation maps that can effectively reuse the data from previous object goals, unlike VLFM, which builds a query-specific scoring map that cannot leverage prior searches. As a result, our method sees a relative increase in efficiency (SPL) by a factor of 1.5x compared to VLFM for the third object, benefiting from the high-quality open-vocabulary semantic map constructed during the search for the first and second objects. We note that $100\%$ SPL here represents perfect (oracle) knowledge of where every object is in advance and is therefore an unattainable upper bound.
@@ V-C Real World Experiments
We deploy OneMap on a Boston Dynamics Spot quadruped using only one front-facing Realsense D455 stereo depth camera, and a Livox Mid 360 lidar with Fast-LIO2 [43] for odometry. We use Yolo-World [39] instead of YOLOv7 to allow searching arbitrary objects, instead of MS-COCO [38] classes only. We run the complete system on-board a Jetson Orin AGX, and achieve an overall update frequency of $2\,\mathrm{Hz}$, which we found to be sufficient for an indoor, legged robot.
@@ VI Conclusion
We found that our method works well with projection to 2D, but future work could include examining suitable 3D map representations. The presented mapping method can readily be adapted to produce a 3D voxel map, though this will introduce additional complexity to the mapping and planning parts. Moreover, means to address noise stemming from the reliance on depth images should be addressed. Lastly, the applicability of the method to outdoor scenarios where semantics might be less meaningful needs to be studied.